### 25 Feb 2024 v1

@@ Abstract
Combining machine learning (ML) with computational fluid dynamics (CFD) opens many possibilities for improving simulations of technical and natural systems. However, CFD+ML algorithms require exchange of data, synchronization, and calculation on heterogeneous hardware, making their implementation for large-scale problems exceptionally challenging.
We provide an effective and scalable solution to developing CFD+ML algorithms using open source software OpenFOAM and SmartSim. SmartSim provides an Orchestrator that significantly simplifies the programming of CFD+ML algorithms and a Redis database that ensures highly scalable data exchange between ML and CFD clients. We show how to leverage SmartSim to effectively couple different segments of OpenFOAM with ML, including pre/post-processing applications, solvers, function objects, and mesh motion solvers. We additionally provide an OpenFOAM sub-module with examples that can be used as starting points for real-world applications in CFD+ML [^1] https://github.com/OFDataCommittee/openfoam-smartsim.
@@ 1 Introduction
Machine learning (ML) and artificial intelligence (AI) methods are increasingly being applied to scientific research, with the field of computational fluid dynamics (CFD) being no exception. This has led to the emergence of at least two hybrid AI/numerical simulation paradigms: AI-in-the-loop and AI-outside-the-loop. Both of these are distinguished by the coupling of an AI method with the simulation to form a new complex CFD+ML algorithm. In the case of AI-in-the-loop, the AI method is embedded within the simulation as a part of the numerical solver. For example, this could be an artificial neural network (ANN) surrogate model for sub-grid-scale physics or an ML model trained in-situ as the simulation progresses. AI-around-the-loop refers to the application of AI methods which interact with the system via its inputs and outputs. Common examples include automated parameter tuning using black-box optimization techniques which call the simulation. In both cases, the most widely used programming language of choice is Python, though common ML frameworks provide C++ Application Programming Interfaces (API) as well.
@@
In response to these challenges, the OpenFOAM Special Interest Group for Data-driven Modelling has been considering the following design hypothesis: scientific simulations should be loosely coupled to ML both from a software sense (e.g., the ML backends should not be linked directly into the application) and also in a compute sense (e.g., ML training and inference should be offloaded to a separate process). This loose coupling allows for a clean separation of concerns and delivers modularity when examining different ML frameworks. Additionally, we consider a data-sharing paradigm which uses a centralized, computation-enabled database to stage data in-memory.
In this paper, we implement this loosely coupled paradigm within the OpenFOAM context using the open-source SmartSim library developed by Hewlett Packard Enterprise. We demonstrate three use cases as examples of these hybrid CFD+ML workflows:
  • Bayesian optimization for tuning the parameters of a turbulence model with the goal of matching the results of a low-resolution model with a higher fidelity one
  • Streaming CFD data from the simulation to calculate its reduced basis via a partitioned singular value decomposition (SVD)
@@
Underlying these use cases is the expression of a new philosophy for scientific computation which considers workflow components as data producers and consumers. We focus on educational examples that can be easily understood and reproduced by the computational science community, and used as implementation starting points by the OpenFOAM community for developing more complex CFD+ML applications. The examples in this paper are therefore minimal working examples of complex CFD+ML workflows, representative of future large-scale applications that are currently in development.
@@ 2.1 Computational and communication architecture
In a loosely coupled CFD+ML framework, both the CFD and ML algorithm can be considered separate entities and, thus, also deployed separately. In this paper, we consider loosely coupled CFD+ML algorithms to also mean that ML and CFD algorithms do not share memory. This creates a need to communicate data between the entities. Anticipating more complex workflows in the future, we reject the concept of a peer-to-peer communication model as the number of connections between entities scales by a power law which becomes untenable at even modest scales. Consider an OpenFOAM case deployed on 1024 ranks, a second computational component deployed on 32 ranks, and a third application on 16 ranks. This configuration leads to a total of 524,288 connections. The problem is further amplified if applications might join or leave the ecosystem at various points of the workflow.
@@
We next demonstrate the additional functionality gained if the distributed database is capable of computation. With this simple addition, data can be sent, transformed, and made available to any entity in the ecosystem. In the context of a hybrid CFD+ML workflow, this database can be used an inference server, and thus, the other entities do not require their own specialized hardware. SmartSim implements this distributed database using a Redis cluster that can be sharded over a number of nodes, as explained in next section. Communication clients, provided by the related SmartRedis library and embedded in the application, connect to this database on application launch. The embedding of these clients into OpenFOAM is discussed in more detail in 2.3.
@@ 2.2 SmartSim and SmartRedis
To enable the CFD+ML use cases described in this work, two libraries are used together with OpenFOAM: SmartSim [1], which sets up and runs the workflow’s computational infrastructure, and SmartRedis, a client to Redis databases used to exchange data across workflow applications and to run Machine Learning (ML) models and other processing functions on such data.
@@
From a software development standpoint, SmartRedis is lightweight and minimally invasive: only a few calls to SmartRedis functions need to be added to existing application code to allow the integration in the orchestrated workflow. SmartRedis can be used to instrument applications written in C, C++, Fortran, and Python. SmartRedis has been tested on parallel applications running at scale on state-of-the-art supercomputers [2], and its impact on the overall workflow computational performances is very low, especially considering the richness of features it adds to standard numerical applications.
@@ 2.3 Communication and signaling
The loose coupling of CFD+ML algorithms using OpenFOAM and SmartSim, enables fast design and deployment to problem-specific data-driven asynchronous and synchronous workflows. Consider, for example, fig. 1, that contains a CFD+ML workflow developed with OpenFOAM and SmartSim for live post-processing of CFD results using a generic ML model. A concrete example of this workflow is described in section 4.2 for a distributed version of the SVD applied to CFD results from OpenFOAM. As shown in section 4.2, an OpenFOAM solver or a function-object communicates OpenFOAM fields to the SmartRedis database. A SmartSim implementation of the ML workflow fetches OpenFOAM fields and trains the ML model to approximate the fields. Both the simulation loop and approximation loop are iterating in user-defined frequencies. The simulation loop may iterate over time steps or pseudo time steps, while the approximation loop may iterate over one or more snapshots of flow fields. The nature of the model defines the signaling frequency between the simulation loop and the approximation loop. The signaling is achieved by storing (and deleting) flags in the SmartSim database. In fig. 1, the CFD algorithm communicates at the end of the simulation an appropriate flag to the ML algorithm through SmartRedis. In a batch-mode ML approximation loop, the signal would notify the ML algorithm that a sufficient number of simulation loop iterations have passed and the time has come for the ML algorithm to approximate batched OpenFOAM fields stored in SmartRedis.
@@
Figure 1: Online Machine-Learning approximation of CFD results from OpenFOAM in SmartSim. Full lines represent transitions between algorithm steps, dashed lines are database operations.
@@
A bidirectional loosely-coupled CFD+ML workflow requires more complex communication and signaling between the CFD and ML algorithms, as shown in fig. 2. Any CFD+ML workflow where the CFD algorithm requires data from the ML algorithm repeatedly after a number of CFD iterations requires bidirectional communication and signaling. An example of a highly challenging bidirectional loosely-coupled CFD+ML workflow, described in detail in section 4.3, is the approximation of displacements of the unstructured finite volume mesh using an artificial neural network (ANN), that is trained on new data every time step in the CFD simulation.
@@
Figure 2 contains a schematic representation of a generic bidirectional CFD+ML workflow, realized with OpenFOAM, SmartSim, and SmartRedis. The CFD algorithm communicates CFD data to the SmartRedis database in the form of SmartRedis tensors. The SmartSim entity, also a client of the SmartRedis database, implements the ML algorithm, which obtains the tensors needed to train the ML model. Once the ML model has been trained on CFD data, it is stored in the SmartRedis database. Concurrently with these operations, each MPI rank of the CFD algorithm (when run in parallel) sends the data for forward inference to SmartRedis and queries the SmartRedis database for an ML model availability flag. Once the ML model availability flag has been set by the ML client of the SmartRedis database, the CFD algorithm requests the forward inference within SmartRedis. Tensors resulting from the forward inference in SmartRedis are obtained by the OpenFOAM CFD client and used in a simulation. As in fig. 1, the bidirectional CFD+ML algorithm from fig. 2 synchronizes the approximation loop and the CFD loop using a flag for the end of the simulation, stored in SmartRedis by the CFD algorithm, and polled for in the SmartRedis database by the SmartSim client.
@@ 3 OpenFOAM and SmartSim integration
The loose CFD+ML coupling described in section 2 significantly simplifies implementing the integration of SmartSim and OpenFOAM. The SmartSim and SmartRedis application programming interfaces (API) are very concise, because loose coupling only requires an orchestration of OpenFOAM simulations as single or multiple, blocking or non-blocking elements of CFD+ML workflows (cf. figs. 1 and 2), sending and receiving tensors to and from SmartRedis, forward inference of ML models in SmartRedis, and writing and polling signaling flags in the SmartRedis database [^2] https://www.craylabs.org/docs/overview.html.
Therefore, as a C++ client of the SmartRedis database, any element the OpenFOAM CFD algorithm that takes part in communication and signaling described in section 2.3 links to the smartredis library, opens a communication channel to SmartRedis, and uses the SmartRedis C++ API [^3] https://www.craylabs.org/docs/api/smartredis_api.html#c to perform above the mentioned tasks.
Contrary to the classical CFD workflow, where OpenFOAM solvers are independently started in simulations, the loosely-coupledwCFD+ML workflow introduces dependencies to OpenFOAM in the form of the database and the ML algorithm. The interaction between OpenFOAM, SmartSim and the ML algorithm is governed by the SmartSim driver script, which starts OpenFOAM simulations as SmartSim Models.
The following section describes a generalization of the communication between OpenFOAM and SmartRedis, by implementing the integration of SmartRedis in an OpenFOAM function object. OpenFOAM function objects are elements of OpenFOAM that enable user-defined calculations within a simulation, which be attached to any OpenFOAM solver at runtime via configuration files, without modifying the solver [3].
@@ 3.1.1 Standardizing OpenFOAM-SmartRedis interactions
This section describes an (API) for interfacing OpenFOAM with the SmartRedis in-memory database. The API is organized into three layers, which support varying levels of control over the interaction with the SmartRedis database.
@@ 3.1.2 A function object for interacting with SmartRedis
The proper usage of such function objects to send pressure, velocity, and face flux data for both the internal field and the inlet patch, is shown in fig. 6, where both field names and target patches can be selected dynamically by the user. A tutorial case showcasing how this function object can be used is provided in [^4] https://github.com/OFDataCommittee/openfoam-smartsim/blob/e72bb32/tutorials/functionObjectthe associated repository.
@@
Fetching the data in the ML code requires knowing the naming convention, echoed back by the function object itself, as shown in fig. 7. The Jinja2 templating engine can be used to hot-replace the placeholders with their desired values in the ML code if needed. To illustrate, the SmartRedis database needs to be queried for the {pUPhi_time_index_0_mpi_rank_0}.field_name_patch_inlet tensor if the user is looking for the inlet data of the pressure field from MPI rank 0 at the first time index.
@@
Another important component of the function object is the metadata dataset that is automatically submitted to the SmartRedis database. This special dataset will be named pUPhi_metadata for the example in fig. 6, and is intended for handling metadata information that the master process wants to communicate to the ML code. These metadata are case-specific, but they will always include the naming conventions for datasets and fields. To illustrate, fig. 8 shows how this metadata dataset can be exploited to derive corresponding tensor names for fields in a generic and automated way through Python scripting. Note that in case SmartSim is running an ensemble-based model for parameter variation, if data is to be fetched from the database for a specific ensemble member, only an environment variable needs to be set (SSKEYIN) in order for the name returned by get_field_name from fig. 8 to be a valid database key.
@@
ds_name = ds_naming.render(time_index=timestep, mpi_rank=processor)

### 23 Apr 2024 v2

@@ Abstract
This is the accepted preprint of the published article: https://doi.org/10.1007/s11012-024-01797-z, please refer to the published article when citing this work.
Combining machine learning (ML) with computational fluid dynamics (CFD) opens many possibilities for improving simulations of technical and natural systems. However, CFD+ML algorithms require exchange of data, synchronization, and calculation on heterogeneous hardware, making their implementation for large-scale problems exceptionally challenging. We provide an effective and scalable solution to developing CFD+ML algorithms using open source software OpenFOAM and SmartSim. SmartSim provides an Orchestrator that significantly simplifies the programming of CFD+ML algorithms enables scalable data exchange between ML and CFD clients. We show how to leverage SmartSim to effectively couple different segments of OpenFOAM with ML, including pre/post-processing applications, function objects, and mesh motion solvers. We additionally provide an OpenFOAM sub-module with examples that can be used as starting points for real-world applications in CFD+ML.
@@ 1 Introduction
Machine learning (ML) and artificial intelligence (AI) methods are increasingly being applied to scientific research, with the field of computational fluid dynamics (CFD) being no exception. This has led to the emergence of at least two hybrid AI/numerical simulation paradigms: AI-in-the-loop and AI-outside-the-loop. Both of these are distinguished by the coupling of an AI method with the simulation to form a new hybrdi, CFD+ML algorithm. In the case of AI-in-the-loop, the AI method is embedded within the simulation as a part of the numerical solver. For example, this could be an artificial neural network (ANN) surrogate model for sub-grid-scale physics or an ML model trained in-situ as the simulation progresses. AI-around-the-loop refers to the application of AI methods which interacts with the system via its inputs and outputs. Common examples include automated parameter tuning using black-box optimization techniques which where the simulation output forms a part of the objective function. In both cases, the most widely used programming language of choice is Python, though common ML frameworks provide C++ Application Programming Interfaces (API) as well.
@@
In response to these challenges, the OpenFOAM Special Interest Group for Data-driven Modelling [1] has been considering the following design hypothesis: scientific simulations should be loosely coupled to ML both from a software sense (e.g., the ML backends should not be linked directly into the application) and also in a compute sense (e.g., ML training and inference should be offloaded to a separate process). This loose coupling allows for a clean separation of concerns and delivers modularity when examining different ML frameworks. Additionally, we consider a data-sharing paradigm which uses a centralized, computation-enabled database to stage data in-memory.
In this paper, we implement this loosely coupled paradigm within the OpenFOAM context using the open-source libraries SmartSim and SmartRedis, developed by Hewlett Packard Enterprise. We utilize the latest releases of OpenFOAM (version 2312 [2]), SmartSim (version 0.6.2. [3]), and SmartRedis (version 0.5.2 [4]), and provide three examples of hybrid CFD+ML workflows, available in a publicly available GitHub repository [5] and as a source code archive on the Zenodo data repository [6]:
  • Bayesian optimization for tuning the parameters of a turbulence model with the goal of matching the results of a low-resolution model with a higher fidelity reference case.
  • Streaming CFD data from the simulation to calculate its reduced basis via a partitioned singular value decomposition (SVD).
@@
Underlying these use cases is the expression of a new philosophy for scientific computation which considers workflow components like the simulation and ML applications as data producers and consumers. We focus on educational examples that can be easily understood and reproduced by the computational science community, and used as implementation starting points by the OpenFOAM community for developing more complex CFD+ML applications. The examples in this paper are therefore minimal working examples of complex CFD+ML workflows, representative of future large-scale applications that are currently in development.
@@ 2.1 Computational and communication architecture
In a loosely coupled CFD+ML framework, both the CFD and ML algorithm can be considered separate entities and, thus, also deployed separately. In this paper, we consider loosely coupled CFD+ML algorithms to also mean that ML and CFD algorithms do not share memory. This creates a need to communicate data between the entities. Anticipating more complex workflows in the future, we reject the concept of a peer-to-peer communication model as the number of connections between entities scales by a power law which becomes untenable at even modest scales. Consider an OpenFOAM case deployed on 1,024 ranks, a second computational component deployed on 32 ranks, and a third application on 16 ranks. This configuration leads to a total of 524,288 connections. The problem is further amplified if applications might join or leave the ecosystem at various points of the workflow.
@@
We next demonstrate the additional functionality gained if the distributed database is capable of computation. With this simple addition, data can be sent, transformed, and made available to any entity in the ecosystem. In the context of a hybrid CFD+ML workflow, this database can be used an inference server, and thus, the other entities do not require their own specialized hardware. SmartSim implements this distributed database using a Redis cluster that can be sharded over a number of nodes, as explained in next section. Communication clients, provided by the related SmartRedis library and embedded in the application, connect to this database on application launch. The embedding of these clients into OpenFOAM is discussed in more detail in section 2.3.
@@ 2.2 SmartSim and SmartRedis
To enable the CFD+ML use cases described in this work, two libraries are used together with OpenFOAM: SmartSim [7, 3], which sets up and runs the workflow’s computational infrastructure, and SmartRedis, a client to Redis databases used to exchange data across workflow applications and to run Machine Learning (ML) models and other processing functions on such data.
@@
From a software development standpoint, SmartRedis is lightweight and minimally invasive: only a few calls to SmartRedis functions need to be added to existing application code to allow the integration in the orchestrated workflow. SmartRedis can be used to instrument applications written in C, C++, Fortran, and Python. SmartSim and SmartRedis have been tested on CFD-like parallel applications running at scale on state-of-the-art supercomputers. Results from these applications and discussions about the performance of these CFD+ML applications can be found in references [7, 8, 9]. These integrations are more easily done when the source code is user-modifiable (as in the case of OpenFOAM) or if the application has defined ways of injecting custom code (e.g. via user defined functions or plugins).
The client/orchestrator architecture provided by SmartSim is uniquely suited for high performance computing systems which have heterogeneous nodes, i.e. some nodes may have accelerators whereas other nodes are traditional CPU-only nodes. If the ML backends were linked directly into OpenFOAM, every node that the AI component of OpenFOAM was run on would expect to have a GPU. By decoupling the ML component and using the Orchestrator as an inference engine and/or an in-memory cache for training applications, GPU resources can be limited to only the portion of the workflow that absolutely require them. Additionally, for CPU-based simulation codes like OpenFOAM, the entire AI workload can be handled by a relatively small amount of GPUs because the AI model sizes tend to be small [7, 9]. By scaling the database to the workload, users on these environments can ensure high utilization of the GPU resources.
@@ 2.3 Communication and signaling
The loose coupling of CFD+ML algorithms using OpenFOAM and SmartSim, enables fast design and deployment to problem-specific data-driven asynchronous and synchronous workflows. Consider, for example, fig. 1, that contains a CFD+ML workflow developed with OpenFOAM and SmartSim for live post-processing of CFD results using a generic ML model. A concrete example of this workflow is described in section 4.3 using a distributed version of the SVD applied to CFD results from OpenFOAM. As shown in section 4.3, an OpenFOAM solver or a function-object communicates OpenFOAM fields to the SmartRedis database. A SmartSim implementation of the ML workflow fetches OpenFOAM fields and trains the ML model to approximate the fields. Both the simulation loop and approximation loop iterate at user-defined frequencies. The simulation loop may iterate over time steps or pseudo time steps, while the approximation loop may iterate over one or more snapshots of flow fields.
The nature of the modelled problem defines the signaling frequency between the simulation loop and the approximation loop. The signaling is achieved by storing, deleting and polling flags in the SmartSim database. In this context, polling information from the database is defined as performing repeated queries in user-defined time intervals and maximal number of attemps, until the polled information becomes available. Polling is necessary because SmartRedis internally manages and balances its input/output operations, that are generally asynchronous. Asynchronicity in this sense means that when a client writes to SmartRedis, it will request a write operation - there is no way for us to know when exactly the data will be actually written. Since each data object in SmartRedis has a name (i.e. a key), other clients can very efficiently check if a key is available, and only fetch data once a key becomes available, i.e. poll the database for the key. This data-centric approach is at the core of our workflow, replacing numerous, complicated and inneficient client-to-client communications on heterogeneous hardware with a database-centric poll/get/put operations. In fig. 1, the CFD algorithm communicates at the end of the simulation an appropriate flag to the ML algorithm through SmartRedis. In a batch-mode ML approximation loop, the signal would notify the ML algorithm that a sufficient number of simulation loop iterations have passed and the time has come for the ML algorithm to approximate batched OpenFOAM fields stored in SmartRedis.
@@
Figure 1: Online ML approximation of CFD results from OpenFOAM in SmartSim. Full lines represent transitions between algorithm steps, dashed lines are database operations.
@@
A bidirectional, loosely-coupled CFD+ML workflow requires more complex communication and signaling between the CFD and ML algorithms, as shown in fig. 2. The approximation of displacements of the unstructured finite volume mesh using an ANN is one such example of this workflow (discussed in section 4.4). In this case, the ANN must be updated on with new data every time step in the CFD simulation, called for inference with fields from the current state of the model, and then retrieved to realize the actual movement of hte mesh.
@@
Figure 2 contains a schematic representation of a generic bidirectional CFD+ML workflow, using objects and concepts from OpenFOAM, SmartSim, and SmartRedis. The CFD algorithm communicates CFD data to the SmartRedis database in the form of SmartRedis tensors. A SmartSim entity implements the ML training algorithm (which also contains a SmartRedis client), obtaining the tensors needed to train the ML model. Once the ML model has been trained on CFD data, it is stored in the SmartRedis database. Concurrently with these operations, each MPI rank of the CFD algorithm (when run in parallel) sends the data for forward inference to SmartRedis and queries the SmartRedis database for an ML model availability flag. Once the ML model availability flag has been set by the ML client of the SmartRedis database, the CFD algorithm requests the forward inference within SmartRedis. Tensors resulting from the forward inference in SmartRedis are obtained by the OpenFOAM CFD client and used in a simulation. As in fig. 1, the bidirectional CFD+ML algorithm from fig. 2 synchronizes the approximation loop and the CFD loop using a flag for the end of the simulation, stored in SmartRedis by the CFD algorithm, and polled for in the SmartRedis database by the SmartSim client.
@@ 3 OpenFOAM and SmartSim integration
The loosely coupled, CFD+ML workflows described in section 2 significantly simplifies implementing the integration of SmartSim and OpenFOAM. The SmartSim and SmartRedis application programming interfaces (API) are very concise, because loose coupling requires an orchestration of OpenFOAM simulations with blocking or non-blocking behavior (cf. figs. 1 and 2), sending and receiving tensors to and from SmartRedis, forward inference of ML models in SmartRedis, and writing and polling signaling flags in the SmartRedis database (more information can be found in the SmartSim documentation [10]). These tasks can be achieved by using the C++ client to communicate with the SmartRedis database. An element of the OpenFOAM CFD algorithm that takes part in communication and signaling described in section 2.3 links to the smartredis library, opens a communication channel to SmartRedis, and uses the SmartRedis C++ API [11] to perform above the mentioned tasks.
Contrary to the classical CFD workflow, where OpenFOAM solvers are independently started in simulations, the loosely-coupled CFD+ML workflow introduces dependencies to OpenFOAM in the form of the database and the ML algorithm. The interaction between OpenFOAM, SmartSim, SmartRedis, and the ML algorithm is governed by the SmartSim driver script implemented in Python, which implements the overall workflow, driving OpenFOAM simulations and ML model training as SmartSim Models.
The following section describes a generalization of the communication between OpenFOAM and SmartRedis, by implementing the integration of SmartRedis in an OpenFOAM function object. OpenFOAM function objects are elements of OpenFOAM that enable user-defined calculations within a simulation, which can be attached to any OpenFOAM solver at runtime via configuration files, without modifying the solver [12].
@@ 3.1.1 Standardizing OpenFOAM-SmartRedis interactions
This section describes an API for interfacing OpenFOAM with the SmartRedis in-memory database. The API is organized into three layers, which support varying levels of control over the interaction with the SmartRedis database.
@@ 3.1.2 A function object for interacting with SmartRedis
The proper usage of such function objects to send pressure, velocity, and face flux data for both the internal field and the inlet patch, is shown in fig. 6, where both field names and target patches can be selected dynamically by the user. A tutorial case showcasing how this function object can be used is provided in the associated repository [^1] https://github.com/OFDataCommittee/openfoam-smartsim/blob/e72bb32/tutorials/functionObject.
@@
Fetching the data in the ML code requires knowing the naming convention, echoed back by the function object itself, as shown in fig. 7. The Jinja2 templating engine [13] can be used to hot-replace the placeholders with their desired values in the ML code if needed. To illustrate, the SmartRedis database needs to be queried for the {pUPhi_time_index_0_mpi_rank_0}.field_name_patch_inlet tensor if the user is looking for the inlet data of the pressure field from MPI rank 0 at the first time index.
@@
Another important component of the function object is the metadata dataset that is automatically submitted to the SmartRedis database. This special dataset will be named pUPhi_metadata for the example in fig. 6, and is intended to handle metadata information that the master process wants to communicate to the ML code. These metadata are case-specific, but they will always include the naming conventions for datasets and fields. To illustrate, fig. 8 shows how this metadata dataset can be exploited to derive corresponding tensor names for fields in a generic and automated way through Python scripting. Note that in case SmartSim is running an ensemble-based model for parameter variation, data can be fetched from the database for a specific ensemble member programatically using the set_data_source method on the SmartRedis client for the name returned by get_field_name from fig. 8 to be a valid database key.
@@
%**** ms.tex Line 375 ****ds_name = ds_naming.render(time_index=timestep, mpi_rank=processor)